OpenStudy (anonymous):

A circle has its center at (-1, 2) and a radius of 3 units. What is the equation of the circle? Answer (x + 1)2 + (y - 2)2 = 9 (x - 1)2 + (y + 2)2 = 3 (x + 1)2 + (y - 2)2 = 3 (x + 1)2 + (y + 2)2 = 9

OpenStudy (anonymous):

thanks

OpenStudy (apoorvk):

no its A... radius gets squared... first option is right!!!

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

\[3^2=9\] on the right

myininaya (myininaya):

sat is right

OpenStudy (anonymous):

also apoorvk

OpenStudy (apoorvk):

and so am i too i guess??? :p

myininaya (myininaya):

oops sorry i didn't see apoorvk's respose and him too lol

myininaya (myininaya):

response *

OpenStudy (anonymous):

Oh so you guys are saying its this :(x + 1)2 + (y - 2)2 = 9 and not this (x + 1)2 + (y - 2)2 = 3

OpenStudy (anonymous):

okay thx